OpenStudy (anonymous):

Which function is greater at the given value? y = 5^x or y = x^5 at x = 2. x^5 5^x Could someone explain this to me so I better understand it?

OpenStudy (anonymous):

Substitute x = 2 into each function, and calculate both of them. The first function: \[y = 5^x = 5^2 = 5 \times 5 = 25\] The second function: \[ y = x^5 = 2^5 = 2 \times 2 \times 2\times 2 \times 2 = 32\] So the second function is greater at given value of "x = 2".

OpenStudy (anonymous):

Thankyou so much im writing down notes on everything so I can learn a little bit better and get the hang of it :)